Section 9

Ordinary Differential Equations and Dynamical Systems

In case of several authors, Invited Speakers are marked with a ^{∗}.

W. de Melo: Rigidity and Renormalization in One Dimensional Dynamical Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 765

L. H. Eliasson: Reducibility and Point Spectrum for Linear Quasi-Periodic Skew-Products . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 779

Shuhei Hayashi: Hyperbolicity, Stability, and the Creation of Homoclinic Points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 789

Michael Herman: Some Open Problems in Dynamical Systems …. II 797
Yuri Kifer: Random Dynamics and its Applications . . . . . . . . . . . . . . . . II 809

Sergei B. Kuksin: Elements of a Qualitative Theory of Hamiltonian PDEs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 819

Krystyna Kuperberg: Counterexamples to the Seifert Conjecture . II 831

Curtis T. McMullen: Rigidity and Inflexibility in Conformal Dynamics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 841

Grzegorz Świa̧tek: Induced Hyperbolicity for One-Dimensional Maps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 857

Zhihong Xia: Arnold Diffusion: A Variational Construction . . . . . . . . . II 867

Documenta Mathematica · Extra Volume ICM 1998 · II ·

764

Documenta Mathematica · Extra Volume ICM 1998 · II ·